OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1

OpenStudy (camerondoherty):

Ok so Step 1: \[2(n - 3) + 5 = 3(n - 1) \] Distribute \[2 \times n-3 \]and \[3 \times n-1 \] Then your equation looks like this: \[2n - 6 + 5 = 3n - 3 \] Combine like terms \[-6 + 5 = -1\] Then you have: \[2n - 1 = 3n - 3 \] Add 1 to both sides: \[2n=3n-2 \] Then subtract 3n from both sides: \[-n = -2 \] \[n=2 \]

OpenStudy (camerondoherty):

Sorry that it took so long >.<

OpenStudy (camerondoherty):

Im a slow typer

OpenStudy (anonymous):

No Problem!! Thank you!!

OpenStudy (camerondoherty):

You're Welcome C: